# Tony Lindeberg

### v2

@@ 2.2 Idealized models for complex cells
with the underlying space-time separable spatio-temporal receptive fields $T_{\varphi^{m},t^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})$ according to (2.1) for $n=0$.
@@
Figure 4: Graphs of the orientation selectivity for the idealized models of (left column) simple cells in terms of first-order directional derivatives of affine Gaussian kernels, (middle column) simple cells in terms of second-order directional derivatives of affine Gaussian kernels and (right column) complex cells in terms of directional quasi quadrature measures that combine the first- and second-order simple cell responses in a Euclidean way for $C_{\varphi}=C_{t}=1/\sqrt{2}$, and shown for different values of the ratio $\kappa$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the degree of orientation selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.)
2.3 Orientation selectivity curves for the idealized receptive field models
In (Lindeberg 2024) the responses of the above purely spatial models of receptive fields are calculated with respect to a static sine wave of the form (see Figure 3)
@@ 2.3 Orientation selectivity curves for the idealized receptive field models
with the velocity vector $(u_{1},u_{2})^{T}$ parallel to the inclination angle $\theta$ of the grating such that $(u_{1},u_{2})^{T}=(u\cos\theta,u\sin\theta)^{T}$.
In summary, the theoretical analysis in (Lindeberg 2024) shows that the resulting orientation selectivity curves for the first-order simple cells, the second-order simple cells and the complex cells, respectively, thereby will be of the forms
@@
with similar angular dependencies within each class for both the purely spatial receptive fields as for the joint spatio-temporal receptive fields.
@@
Figure 5: Measurements of the orientation tuning of neurons, at different positions in the visual cortex, according to Nauhaus et al. (2008) (Copyright Cell Press with permission). In this figure it can be seen how the orientation tuning changes from broad to sharp, and thus higher degree of orientation selectivity, with increasing distance from the pinwheels, consistent with the qualitative behaviour that would be obtained if the ratio $\kappa$, between the scale parameters in underlying affine Gaussian smoothing step in the idealized models of spatial and spatio-temporal receptive fields, would increase when moving away from the centers of the pinwheels on the cortical surface.
3 Interpretation of the connection between the orientation selectivity and the elongation of receptive fields in relation to biological measurements
@@
In this section, we will compare the results of the above theoretical predictions with biological results concerning the orientation selectivity of visual neurons.
@@
In view of our theoretical results, as summarized in Section 2, concerning the orientation selectivity of receptive fields, where the spatial smoothing part is performed based on affine Gaussian kernels, this qualitative behaviour is consistent with what would be the result if the ratio $\kappa$ between the two scale parameters of the underlying affine Gaussian kernels would increase from a lower to a higher value, when moving away from the centers of the pinwheels on the cortical surface. Thus, the presented theory leads to a prediction about a variability in the eccentricity or the elongation of the receptive fields in the primary visual cortex. In the case of pinwheel structures, the behaviour is specifically consistent with a variability in the eccentricity or the elongation of the receptive fields from the centers of the pinwheels towards the periphery.
Let us furthermore consider the theoretical prediction from (Lindeberg 2023), that the shapes of the affine Gaussian derivative-based receptive fields ought to comprise a variability over a larger part of the affine group than mere rotations, to enable affine covariance and (partial) affine invariance at higher levels in the visual hierarchy. Then, if combined with the theoretical orientation selectivity analysis presented in the paper, those predictions are also consistent with the results by Nauhaus et al. (2008), with an additional explanatory power: If the theoretically motivated prediction would hold, then the underlying theoretical model may enable a deeper interpretation of those biological results, in terms of underlying computational mechanisms in the visual receptive fields, to enable specific functional covariance and invariance properties at higher levels in the visual hierarchy.
A highly interesting quantitative measurement to perform, in view of these theoretical results, would hence be to fit parameterized models of the orientation selectivity, according to Equations (12), (13) and (14)
@@
In (Lindeberg 2023), a theoretical treatment is given concerning covariance properties of visual receptive fields under natural image transformations, specifically geometric image transformations in terms of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations. According to that theory of spatial and spatio-temporal receptive fields, in terms of generalized Gaussian derivatives, the covariance properties of the receptive fields mean that the shapes of the receptive field families should span the degrees of freedom generated by the geometric image transformations. With regard to spatial affine transformations, which beyond spatial scaling transformations do also comprise spatial rotations and non-uniform scaling transformation with different amount of scaling in two orthogonal spatial directions, this theory implies that affine Gaussian kernels ought to be present in the receptive field families corresponding to different values of the ratio $\kappa$ between the spatial scale parameters, to support affine covariance. In (Lindeberg 2023 Section 3.2) suggestions to new biological measurements were further proposed to support (or reject) those hypotheses.
@@
If we would assume that it would be unlikely for the receptive fields to have as strong variability in their orientational selectivity properties as function of the positions of the neurons in relation to the pinwheel structure, as reported in this study, without also having a strong variability in their eccentricity, then by combining the theoretical analysis in this article with the biological results by Nauhaus et al. (2008), that would serve as possible indirect support for the hypothesis concerning an expansion of receptive field shapes over variations in the ratio between the two scale parameters of spatially anisotropic receptive fields. Thus, if we would assume that the biological receptive fields can be well modelled by the generalized Gaussian derivative model based on affine Gaussian receptive fields, then the biological results by Nauhaus et al. (2008) are fully consistent with the prediction of such an explicit expansion over shapes of the visual receptive fields, based on the orientation selectivity of visual receptive fields, whose spatial smoothing component can be well modelled by affine Gaussian kernels.

### v7

@@ 2.2.2 Idealized models for complex cells
with the underlying space-time separable spatio-temporal receptive fields $T_{\varphi^{m},t^{n},\scriptsize\mbox{norm}}(x_{1},x_{2},t;\;\sigma_{\varphi},\sigma_{t},v,\Sigma_{\varphi})$ according to (2.2.1) for $n=0$.
@@
Figure 4: Graphs of the orientation selectivity for the idealized models of (left column) simple cells in terms of first-order directional derivatives of affine Gaussian kernels, (middle column) simple cells in terms of second-order directional derivatives of affine Gaussian kernels and (right column) complex cells in terms of directional quasi-quadrature measures that combine the first- and second-order simple cell responses in a Euclidean way for $C_{\varphi}=C_{t}=1/\sqrt{2}$, and shown for different values of the ratio $\kappa$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the degree of orientation selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.)
Third-order simple cell Fourth-order simple cell
$\kappa=1$ Refer to caption Refer to caption
$\kappa=2$ Refer to caption Refer to caption
$\kappa=4$ Refer to caption Refer to caption
$\kappa=8$ Refer to caption Refer to caption
Figure 5: Graphs of the orientation selectivity for the idealized models of (left column) simple cells in terms of third-order directional derivatives of affine Gaussian kernels and (right column) simple cells in terms of fourth-order directional derivatives of affine Gaussian kernels and shown for different values of the ratio $\kappa$ between the spatial scale parameters in the vertical vs. the horizontal directions. Observe how the degree of orientation selectivity varies strongly depending on the eccentricity $\epsilon=1/\kappa$ of the receptive fields. (top row) Results for $\kappa=1$. (second row) Results for $\kappa=2$. (third row) Results for $\kappa=4$. (bottom row) Results for $\kappa=8$. (Horizontal axes: orientation $\theta\in[-\pi/2,\pi/2]$. Vertical axes: Amplitude of the receptive field response relative to the maximum response obtained for $\theta=0$.)
2.2.3 Orientation selectivity curves for the idealized receptive field models
In Lindeberg (2024a), the responses of the above purely spatial models of receptive fields are calculated with respect to a static sine wave of the form (see Figure 3)
@@ 2.2.3 Orientation selectivity curves for the idealized receptive field models
with the velocity vector $(u_{1},u_{2})^{T}$ parallel to the inclination angle $\theta$ of the grating, such that $(u_{1},u_{2})^{T}=(u\cos\theta,u\sin\theta)^{T}$.
In summary, the theoretical analysis in Lindeberg (2024a) shows that the resulting orientation selectivity curves for the first-order simple cells, second-order simple cells and complex cells, respectively, will be of the forms:
@@
with similar angular dependencies within each class for both the purely spatial receptive fields and the joint spatio-temporal receptive fields.
@@
In Appendix A.1, we additionally extend these results to orientation selectivity for purely spatial third-order simple cells and fourth-order simple cells of the forms
$\begin{split}r_{\scriptsize\mbox{simple},3}(\theta)&=\frac{\left|\cos\theta\right|^{3}}{(\cos^{2}\theta+\kappa^{2}\sin^{2}\theta)^{3/2}},\end{split}$ (19)
$\begin{split}r_{\scriptsize\mbox{simple},4}(\theta)&=\frac{\cos^{4}\theta}{(\cos^{2}\theta+\kappa^{2}\sin^{2}\theta)^{2}},\end{split}$ (20)
and with examples of graphs of these curves shown in Figure 5 for a few values of the scale ratio parameter $\kappa$.
As can be seen from a comparison with the orientation selectivity curves for the idealized models of first-order and second-order simple cells in Figure 4, the orientation selectivity curves of the idealized models of third-order and fourth-order simple cells in Figure 5 follow the trends for the first-order and second-order simple cells in Figure 4, in that the orientation selectivity curves become more narrow, both with increasing values of the scale parameter ratio $\kappa$ and with increasing order of spatial differentiation.
2.3 Relations to previous work
Beyond the works by Nauhaus et al. (2008) and by Goris et al. (2015), that the treatment in Section 3 will largely build upon, there is a large body of work on characterizing the orientation selectivity of neurons, by Watkins and Berkley (1974), Rose and Blakemore (1974), Schiller et al. (1976), Albright (1984), Ringach et al. (2002), Nauhaus et al. (2008), Scholl et al. (2013), Sadeh and Rotter (2014) and Sasaki et al. (2015), as well as concerning biological mechanisms for achieving orientation selectivity by Somers et al. (1995), Sompolinsky and Shapley (1997), Carandini and Ringach (1997), Lampl et al. (2001), Ferster and Miller (2000), Shapley et al. (2003), Seriès et al. (2004), Hansel and van Vreeswijk (2012), Moldakarimov et al. (2014), Gonzalo Cogno and Mato (2015), Priebe (2016), Pattadkal et al. (2018), Nguyen and Freeman (2019), Merkt et al. (2019), Wei et al. (2022) and Wang et al. (2024). The focus of this paper, however, is not on the neural mechanisms that lead to orientation selectivity, but on purely functional properties at the macroscopic level.
Mathematical models of biological receptive fields have beyond in terms of Gaussian derivatives (Koenderink and van Doorn 1984; 1987; 1992; Young and his co-workers 1987; 2001; 2001; Lindeberg 2013a; 2021) also been formulated in terms of Gabor filters (Marcelja 1980; Jones and Palmer 1987a; 1987b; Porat and Zeevi 1988). Gaussian derivatives have, in turn, been used as primitives in theoretical models of early visual processing by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
Hubel and Wiesel (1959; 1962; 1968; 2005) pioneered the study of simple and complex cells. The properties of simple cells have been further characterized by DeAngelis et al. (1995; 2004), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008), Walker et al. (2019) and De and Horwitz (2021), and the properties of complex cells investigated by Movshon et al. (1978), Emerson et al. (1987), Martinez and Alonso (2001), Touryan et al. (2002; 2005), Rust et al. (2005), van Kleef et al. (2010), Goris et al. (2015), Li et al. (2015) and Almasi et al. (2020), as well as modelled computationally by Adelson and Bergen (1985), Heeger (1992), Serre and Riesenhuber (2004), Einhäuser et al. (2002), Kording et al. (2004), Merolla and Boahen (2004), Berkes and Wiscott (2005), Carandini (2006), Hansard and Horaud (2011), Franciosini et al. (2019), Lindeberg (2020), Lian et al. (2021), Oleskiw et al. (2024) and Yedjour and Yedjour (2024). In this work, we have followed a specific way of modelling simple and complex cells in terms of affine Gaussian derivatives, according to the generalized affine Gaussian derivative model for visual receptive fields.
Properties of cortical maps in the primary visual cortex have, in turn, been studied in detail by Bonhoeffer and Grinvald (1991), Blasdel (1992), Maldonado et al. 1997, Koch et al. (2016), Kremkow et al. (2016), Najafian et al. (2022), Jung et al. (2022), Fang et al. (2022) and Vita et al. (2024).
There have been some previous mentions regarding receptive fields with different aspect ratios (Tinsley et al. 2003, Xu et al. 2016), but according to our not knowledge not any in-depth studies in relation to a systematic expansions of receptive field shapes over the degree of elongation, nor any previously reported connections between such expansions of receptive field shapes over the degree of elongation in relation to the property of affine covariance for the family of visual receptive fields.
3 Results
@@
In this section, we will compare the results of the theoretical predictions in Section 2.2 with biological results concerning the orientation selectivity of visual neurons.
For later purposes, we will, however, first extend the above results concerning the orientation selectivity curves for idealized models of simple cells according to the generalized Gaussian derivative model for visual receptive fields, from first-order and second-order simple cells to also comprise third-order and fourth-order simple cells.
The results derived in Section A.1 will then be used in Section 3.1.3, when extending the interpretation in the following Section 3.1.2, based on spatial derivatives up to order 2, to spatial derivatives up to order 4. Readers, who are more interested in the functional results of the theory than the details of the mathematical derivations, should be able to, without major loss of continuity, skip the details in Section A.1, while noting the result summary in Section A.1.3, to then continue directly with Section 3.1.
Figure 6: Measurements of the orientation tuning of neurons, at different positions in the visual cortex, according to Nauhaus et al. (2008) (Copyright Cell Press with permission). From these experimental results, it can be seen how the orientation tuning changes from broad to sharp, and thus higher degree of orientation selectivity, with increasing distance from the pinwheels, consistent with the qualitative behaviour that would be obtained if the ratio $\kappa$, between the scale parameters in underlying affine Gaussian smoothing step in the idealized models of spatial and spatio-temporal receptive fields, would increase when moving away from the centers of the pinwheels on the cortical surface.
3.1 Interpretation of the connection between the orientation selectivity and the elongation of receptive fields in relation to biological measurements
3.1.1 Interpretation of the measurements about broad vs. sharp orientation selectivity of neurons by Nauhaus et al. (2008)
@@
Thus, first of all, we can see that the variability in the orientation selectivity of visual neurons is consistent with a variability in the degree of elongation of the receptive fields, as predicted by the underlying normative theory of visual receptive fields, and specifically the more general hypotheses, that the family of receptive fields ought to be expanded over the degrees of freedom of affine image transformations to be affine covariant. As previously described in the introduction, such affine covariance would support more accurate performance of the visual perception system, as the visual stimuli undergo the variabilities generated by observing objects in the world from different viewing directions. The experimental results by Nauhaus et al. (2008) are from this viewpoint consistent with the receptive field shapes being expanded over different degrees of elongation.
Secondly, in view of our theoretical results, as summarized in Section 2.2, concerning the orientation selectivity of receptive fields, where the spatial smoothing part is performed based on affine Gaussian kernels, this qualitative behaviour is consistent with what would be the result if the ratio $\kappa$ between the two scale parameters of the underlying affine Gaussian kernels would increase from a lower to a higher value, when moving away from the centers of the pinwheels on the cortical surface. Thus, by combination with the experimental results by Nauhaus et al. (2008), the presented theory leads to a prediction about a systematic variability in the eccentricity or the elongation of the receptive fields in the primary visual cortex. In the case of pinwheel structures, such a behaviour is specifically consistent with a variability in the eccentricity or the elongation of the receptive fields from the centers of the pinwheels towards the periphery.
Thus, considering the theoretical prediction from Lindeberg (2023), that the shapes of the affine Gaussian derivative-based receptive fields ought to comprise a variability over a larger part of the affine group than mere rotations, to enable affine covariance and (partial) affine invariance at higher levels in the visual hierarchy. Then, if combined with the theoretical orientation selectivity analysis presented in the paper, those predictions are also consistent with the results by Nauhaus et al. (2008), with an additional explanatory power: If the theoretically motivated prediction would hold, then the underlying theoretical model may enable a deeper interpretation of those biological results, in terms of underlying computational mechanisms in the visual receptive fields, to enable specific functional covariance and invariance properties at higher levels in the visual hierarchy.
A highly interesting quantitative measurement to perform, in view of these theoretical results, would hence be to fit parameterized models of the orientation selectivity, according to Equations (16), (17), (18), (46) and (47)
@@
In Lindeberg (2023; 2024b), theoretical treatments are given concerning covariance properties of visual receptive fields under natural image transformations, specifically geometric image transformations in terms of spatial scaling transformations, spatial affine transformations, Galilean transformations and temporal scaling transformations. According to that theory of spatial and spatio-temporal receptive fields, in terms of generalized Gaussian derivatives, the covariance properties of the receptive fields mean that the shapes of the receptive field families should span the degrees of freedom generated by the geometric image transformations. With regard to spatial affine transformations, which beyond spatial scaling transformations do also comprise spatial rotations and non-uniform scaling transformation with different amount of scaling in two orthogonal spatial directions, this theory implies that affine Gaussian kernels ought to be present in the receptive field families corresponding to different values of the ratio $\kappa$ between the spatial scale parameters, to support affine covariance. In Lindeberg (2023 Section 3.2) suggestions for new biological measurements were further proposed to support (or reject) those hypotheses.
@@
If we would assume that it would be unlikely for the receptive fields to have as strong variability in their orientational selectivity properties as a function of the positions of the neurons in relation to the pinwheel structure, as reported in this study, without also having a strong variability in their eccentricity, then by combining the theoretical analysis in this article with the biological results by Nauhaus et al. (2008), that would serve as possible indirect support for the hypothesis concerning an expansion of receptive field shapes over variations in the ratio between the two scale parameters of spatially anisotropic receptive fields. Thus, if we would assume that the biological receptive fields can be well modelled by the generalized Gaussian derivative model based on affine Gaussian receptive fields, then the biological results by Nauhaus et al. (2008) are fully consistent with the prediction of such an explicit expansion over shapes of the visual receptive fields, based on the orientation selectivity of visual receptive fields, whose spatial smoothing component can be well modelled by affine Gaussian kernels.
Figure 7: Distributions of the absolute value of the resultant $|R|\in[0,1]$ according to (22) for the directional selectivity of visual neurons over populations of simple cells and complex cells, respectively, in the primary visual cortex, from neurophysiological recordings of Macaque monkeys by Goris et al. (2015).
3.1.2 Interpretation of the measurements of orientation selectivity histograms by Goris et al. (2015) based on spatial derivatives up to order 2
These predictions are furthermore consistent with existing biological results by Goris et al. (2015), concerning the distribution of the degree of orientation selectivity of the neurons in the primary visual cortex. By measuring the absolute value $|R|$ of the complex-valued resultant, given by
[…]
for the orientation selectivity curve for each visual neuron, and then computing a normalized histogram of these measurements (see Figure 7), Goris et al. (2015) demonstrate a substantial variability in the orientation selectivity of the receptive fields of simple cells and complex cells in the primary visual cortex.
$R_{\scriptsize\mbox{complex}}=\frac{\kappa^{2}\left(48\left(\kappa^{2}-1\right)^{3/4}\Gamma\left(\frac{5}{4}\right)\,_{2}F_{1}\left(\frac{1}{2},1;\frac{3}{4};\frac{1}{\kappa^{2}}\right)-16\left(\kappa^{2}-1\right)^{3/4}\Gamma\left(\frac{5}{4}\right)+3\sqrt{2\pi}\,\kappa\,\Gamma\left(-\frac{1}{4}\right)\right)}{2\left(\kappa^{2}-1\right)\left(16\left(\kappa^{2}-1\right)^{3/4}\Gamma\left(\frac{5}{4}\right)\,_{2}F_{1}\left(\frac{1}{2},1;\frac{3}{4};\frac{1}{\kappa^{2}}\right)+\sqrt{2\pi}\,\kappa\,\Gamma\left(-\frac{1}{4}\right)\right)}$ (23)
[…]
First-order simple cell Second-order simple cell Complex cell
Refer to caption Refer to caption Refer to caption
Third-order simple cell Fourth-order simple cell
Refer to caption Refer to caption
Figure 9: Graphs of the resultant $R$ for idealized models of (top left) a first-order simple cell according to (LABEL:eq-resultant-1st-order-simple), (top middle) a second-order simple cell according to (LABEL:eq-resultant-2nd-order-simple), (top right) a complex cell according to (26) and Figure 8, (bottom left) a third-order simple cell according to (LABEL:eq-resultant-3rd-order-simple) and (bottom middle) a fourth-order simple cell according to (LABEL:eq-resultant-4th-order-simple) on a log-linear scale, with the horizontal axis parameterized in terms of the logarithm $K=\log\kappa$ of the scale parameter ratio $\kappa$ in the affine Gaussian derivative model of visual receptive fields. (Horizontal axes: scale parameter ratio $\kappa$. Vertical axes: resultant $R$.)
This consistency can be demonstrated by computing the closed-form expression for the absolute value of the resultant of the orientation selectivity curves according to Equations (16)–(18), as done in (Lindeberg 2024a Section 5.1) for the first-order idealized models of simple cells
[…]
$\begin{split}&=\frac{\kappa\left(\kappa\cosh^{-1}\kappa-\sqrt{\kappa^{2}-1}\right)}{\left(\kappa^{2}-1\right)\cosh^{-1}\kappa},\end{split}$ (24)
for the second-order idealized models of simple cells
[…]
$\begin{split}&=\frac{\kappa}{\kappa+1}.\end{split}$ (25)
as well as for the idealized models of complex cells
[…]
with the explicit expression for the last result in Figure 8.
Let us additionally reparameterize these curves in terms of a logarithmic parameterization $K=\log\kappa$ of the scale parameter ratio $\kappa$, which leads to orientation selectivity curves of the forms shown in the top row in Figure 9. Then, we can see that the experimentally obtained distributions in Figure 7 appear to be reasonably consistent with the assumption of a rather uniform distribution over the logarithmically parameterized scale parameter ratio $K=\log\kappa$.
Such a parameterization would specifically constitute a canonical parameterization, if one would simplify222 More generally, one could instead conceive a uniform joint distribution on a hemisphere, as conceived in Figure 8 in (Lindeberg 2021), which regarding first-order spatial derivatives then leads to a distribution of spatial receptive field shapes of the form shown in Figure 13, and possibly complemented with additional priors to account for how important different local surface orientations in the environment would be for the perceptual process, as well as densely the space of combined image orientations $\varphi$ and scale parameter ratios $\kappa$ would need to be sampled, to support sufficiently good approximations of covariance over that submanifold for the local image measurements performed by the family of spatial receptive fields. In this treatment, we do, however, simplify this problem, by instead considering a uniform distribution over a logarithmic transformation of the scale parameter ratio $\kappa$, which is also easier to handle in closed form calculations, and which may be regarded as a coarse approximation, to compensate for gross phenomena with regard to a non-uniform distribution of receptive field shapes over the scale parameter ratio $\kappa$. the 2-D joint distribution of receptive field shapes over the scale parameter ratio $\kappa$ and the orientation $\varphi$ into two independent 1-D distributions over the scale parameter ratio $\kappa$ and the orientation $\varphi$, respectively, in the idealized model of visual receptive fields according to the generalized Gaussian derivative framework.
Thus, also these biological results are consistent with the working hypothesis about an expansion over the degree of elongation of the receptive fields in the primary visual cortex, as would be implied from the assumption of a family of affine covariant visual receptive fields.
If we would aim at more detailed modelling of the experimentally recorded histograms of the resultant measure of the orientation selectivity curves, as reported by Goris et al. (2015) and as schematically reproduced in Figure 7, we need to consider that there would be more free parameters in the modelling to determine, based on the following arguments:
  • One basic question concerns what range of values of the scale parameter ratio $\kappa$ would be spanned by the receptive fields in the primary visual cortex.
    In the graphs shown in the top row Figure 9, we have used a range of the scale parameter ratio $\kappa$ over a factor 10 from the unit value 1, which results in a maximum value of $R$ for the idealized first-order model of a simple cell of about 0.67. In contrast, for the idealized second-order model of a simple cell, $R$ reaches a maximum value of $R$ about 0.91, while for the complex cell, $R$ assumes a maximum value of about 0.83.
    If this range of scale parameter ratios would be reduced to a lower span, then the range of the possible values of $R$ would be reduced, while the range would be expanded if a wider range of scale ratios $\kappa$ would be implemented.
  • Another basic question concerns the distribution of receptive fields with respect to the order of spatial differentiation.
    In our theoretical analysis so far, we have modelled the simple cells in terms of first- and second-order directional derivatives. Experimental results by Young (1985; 1987) do, however, indicate that receptive fields that can be modelled by Gaussian derivatives up to order 4 may be present in the primary visual cortex.
    For such receptive fields of higher order, we will obtain additionally sharper orientation selectivity curves, which will then assume values of $R$ closer to 1 than for the first- and second-order models of simple cells.
    To reproduce an idealized model of a histogram of the resultant $R$ for a population of simple cells, as shown in the top part of Figure 7, we would therefore have to a assume a distribution of receptive fields over different orders of spatial differentiation.
    With ample reservations for to the possibility of substantial statistical fluctuations in the histograms, as could be influenced by the selection of the actual visual neurons from which the histograms were computed in the experiments, after noting that the histogram accumulated over the simple cells show much larger variability between the bins than the histogram accumulated over the complex cells, one may then ask if the peak in the bins number 2 and 3 from the left in the histogram for simple cells in Figure 7 could be influenced by contributions from simple cells whose spatial components of the receptive fields can be well modelled by first-order derivatives (who span a much lower range of the values of $R$ than derivatives of higher order), as well as if the heavy part in the right part of the histogram, especially around bin number 8 from the left, could be influenced by simple cells that may be well modelled by directional derivatives for higher orders than 1.
In the next section, we will consider idealized model of receptive fields corresponding to higher orders of spatial differentiation, to investigate these matters in more detail.
First-order simple cells Second-order simple cells Complex cells
Refer to caption Refer to caption Refer to caption
Third-order simple cells Fourth-order simple cells
Refer to caption Refer to caption
Figure 10: Examples of histograms of the resultant $R$ over populations of (top left) first-order simple cells, (top middle) second-order simple cells, (top right) complex cells, (bottom left) third-order simple cells and (bottom right) fourth-order simple cells, accumulated over a uniform logarithmic distribution of the scale parameter ratio $\kappa$ over the interval $\kappa\in[1/\kappa_{\max},\kappa_{\max}]$ for $\kappa_{\max}=8$. (Horizontal axes: bin over the resultant $R$. Vertical axes: number of receptive fields in this bin in a discrete simulation.)
Simple cells up to order 2 Simple cells up to order 4
Refer to caption Refer to caption
Figure 11: Examples of combined histograms of the resultant $R$ over populations of simple cells of different order (left) up to order 2 and (right) up to order 4, accumulated over a uniform logarithmic distribution of the scale parameter ratio $\kappa$ over the interval $\kappa\in[1/\kappa_{\max},\kappa_{\max}]$ for $\kappa_{\max}=8$. (Horizontal axes: bin over the resultant $R$. Vertical axes: number of receptive fields in this bin in a discrete simulation.)
3.1.3 Interpretation of the measurements of orientation selectivity histograms by Goris et al. (2015) based on spatial derivatives up to order 4
Let us first compute the resultants for the orientation selectivity curves according to Equations (46) and (47) for the third-order idealized models of simple cells
[…]
$\begin{split}&=\frac{\kappa\left(\sqrt{\kappa^{2}-1}\left(\kappa^{2}+2\right)-3\kappa\cosh^{-1}(\kappa)\right)}{\left(\kappa^{2}-1\right)\left(\kappa\sqrt{\kappa^{2}-1}-\cosh^{-1}(\kappa)\right)},\end{split}$ (27)
and for the fourth-order idealized models of simple cells
[…]
$\begin{split}&=\frac{\kappa\,(\kappa+3)}{(\kappa+1)(\kappa+2)}.\end{split}$ (28)
From the graphs of these resultant curves as function of the logarithm of the scale parameter ratio $\kappa$ shown in the bottom row in Figure 9, we can with comparison to the graphs in the top row in Figure 9 note that the distributions of resultant $R$ are heavier towards larger values $R$ for the third-order and fourth-order models of simple cells than for the second-order simple cells or the complex cell model based on first-order and second-order directional derivatives. The maximum values of $R$ for the third-order and fourth-order models of simple cells are also higher than for the first-order and second-order models, with the maximum value being about 0.97 for the third-order model and about 0.98 for the fourth-order model.
From such a viewpoint, it seems plausible that a distribution of receptive fields over range of values of the scale parameter ratio $\kappa$ as well as over different orders of spatial differentiation could lead to a bump in the histogram of the resultant $R$ for somewhat larger values of $R$, as can be seen in the experimental results for simple cells reported by Goris et al. (2015) and as reproduced in Figure 7.
Let us next compute the histograms over the resultant $R$ that will be the result if we would assume that the scale parameter ratio $\kappa$ would be uniformly distributed on a logarithmic scale $K=\log\kappa$ over some range delimited by $\kappa_{\min}$ and $\kappa_{\max}$. As previously remarked in Footnote 2, such an assumption may constitute a simplification, since one may argue that a uniform distribution on a hemisphere, complemented with possible additional priors to take into account the accuracy of sampling in the parameter space, could be more appropriate. The assumption of a logarithmic distribution does, however, constitute a principled first assumption to use when to parameterize a strictly positive variable, that may assume values over an extended interval on the positive axis, in terms of natural coordinates, in the absence of further information (Jaynes 1968).
To illustrate to what extent the distributions of the resultant will be influenced by receptive fields for different orders of spatial differentiation, the top row in Figure 10 shows histograms of the resultant accumulated for idealized models of simple cells of orders 1 and 2 as well as for the idealized model of complex cells, based on a quasi-quadrature combination of the responses from first-order and second-order simple cells. For generating these graphs, we have created a uniform distribution of the scale parameter ratio $\kappa$ over a logarithmic scale, over the interval $\kappa\in[1/\kappa_{\max},\kappa_{\max}]$ for the arbitrary choice of $\kappa_{\max}=8$. The bottom row in Figure 10 shows corresponding histograms of the resultant for third-order and fourth-order simple cells.
As can be seen from these graphs, the histogram over the first-order simple cells is delimited by a maximum value around $R_{\max,1}\approx 0.7$, while the distributions for third-order and fourth-order simple cells are heavier for larger values of the resultant approaching $R\rightarrow 1$, and in line with the above arguments. The model used for computing histograms of the resultant for the idealized models of complex cells does, however, not very well reproduce the shape of the biologically obtained histogram, thus indicating that the model for the complex cells may be overly simplified [^3] In this context, it should be remarked that there is strong conceptual difference between the idealized models for simple cells vs. the idealized models of complex cells. The idealized model for simple cells has been determined in a theoretically principled manner from axiomatic derivations, and also been matched to biological measurements of simple cells, whereas the idealized model for complex cells has been chosen as an as straightforward way as possible for combining the responses of odd-shaped and even-shaped simple cells. for the purpose of reproducing the shape of the resultant histogram, see the discussion in Section 6 in (Lindeberg 2024a) for a number of suggestions concerning ways to extend that model.
Figure 11 shows additional results of combining the resultant for populations of simple cells over different orders of spatial differentiation, either up to order 2 or up to order 4, here assuming the same number of neurons for all the different orders of spatial differentiation.
With ample reservation from the facts that: (i) in the current state, we have not principled biological arguments for choosing particular values of the parameters $\kappa_{\min}$ and $\kappa_{\max}$ that determine the range of the scale parameter ratio $\kappa$, where different choices of these parameters may affect the shapes of the combined histograms of the resultant $R$, (ii) the choice of a logarithmic distribution over the scale parameter $\kappa$ does, as previously mentioned, neglect any co-dependency with respect to the distribution over the orientation angle $\varphi$, (iv) the assumption about equal numbers of receptive fields for the different orders of spatial differentiation may not necessarily hold in reality, and additionally expressing reservation from the fact that (v) our computations of the resultant $R$ for the receptive fields are based on idealized noise free model, while there additionally could be sources to noise in the biological experiments as well as modelling errors between the receptive fields of the actual biological neurons and our idealized receptive fields, as can be seen from these results, these combined histograms give rise to a bump in the histograms for lower values of the resultant $R$, in qualitative agreement with the biological results by Goris et al. (2015) and as reproduced in Figure 7. Furthermore, to obtain something that would look like a small bump for larger values of the resultant $R$, the modelling situation with receptive fields up to order 4 would give a closer similarity to the biological results by Goris et al. (2015) compared to the model based on receptive fields up to order 2.
From this analysis, it thus appears as if (i) the histogram of the resultant of the simple cells in the biological experiments et al. (2015) could be rather well explained by the receptive fields in the primary visual cortex of Macaque monkeys having a variability over the degree of elongation, and that additionally (ii) the non-uniform nature of the experimentally obtained histogram of the resultant $R$ for the simple cells could be explained better by assuming that receptive fields should be present up to a spatial differentiation order up to 4 than up to a spatial differentiation order of 2.
3.1.4 Summary of the interpretations of the biological experiments by Nauhaus et al. (2008) and by Goris et al. (2015)
With regard to the working hypothesis that we set out to investigate the possible validity of, we can conclude that the experimental results by both Nauhaus et al. (2008) and by Goris et al. (2015) are, in combination with the theoretical results in Sections 2.2 and A.1, concerning a direct connection between the degree of elongation of a receptive field with the orientation selectivity of the receptive field, clearly consistent with an expansion over the degree of elongation of the receptive field shapes in the primary visual cortex.